OpenStudy (anonymous):

Suppose f(x) is a differentiable function with f(1)=2, f(2)=-2, f'(2)=3,f'(1)=5 and f(5)=1 State an equation of a line is a tangent to the graph of f.

OpenStudy (anonymous):

Assume the tangent line is y' = mx + b then 5 = m + b 3 = 2m + b so m = -2 and b = 7. y' = -2x + 7

OpenStudy (anonymous):

Did you have to use f'(1)=5 to find the answer?

OpenStudy (anonymous):

5 = m + b uses f'(1)=5 3 = 2m + b uses f'(2)=3 I assume there is other parts to the question that u need the other data for (eg u can find the curve).

OpenStudy (anonymous):

Hmm this is the whole question.There is no other part. So for f'(1)=5 is using coordinate (1,5) Then for f(1)=2 is using coordinate (2,1)?

OpenStudy (anonymous):

So for f'(1)=5 is using coordinate (1,5)-- I used this one. Then for f(1)=2 is using coordinate (2,1)? I didn't use this one, if you are going to use it for something, f(1) means "the value of the function when x is 1 ie the x value) and 2 is the result or y value.

OpenStudy (anonymous):

What is the difference for both of this? since both y value are positioned at the right side of the equation.

OpenStudy (anonymous):

I don't understand your question. The difference of what?

OpenStudy (anonymous):

f(x) and f'(x)

OpenStudy (anonymous):

The first one means "function of x" and the second one is another way of writing dy/dx (the derivative).

OpenStudy (anonymous):

So how to do you know f'(1)=5 is using the cordinate (1,5) instead of (5,1)?

OpenStudy (anonymous):

Same like the other, it means the value of the derivative when x is 1 is 5 (ie the y value).

OpenStudy (anonymous):

Let me show u in a more concrete way....

OpenStudy (anonymous):

Oh i see.Thanks !! btw do you know what is the difference between f'(x) and f\[^{-1}\](x) ?

OpenStudy (anonymous):

[f ^{-1}(x)\]

OpenStudy (anonymous):

sorry typo. \[f^{-1}(x)\]

OpenStudy (anonymous):

Does that clarify things a bit?

OpenStudy (anonymous):

Hmm yea a little.thanks!

OpenStudy (anonymous):

f^-1 (f(x)) = x It is the inverse function, which reverses the action of a function.

OpenStudy (anonymous):

hmm thanks!! may i know how you draw the graph?

OpenStudy (anonymous):

Paint accessory.

OpenStudy (anonymous):

cool thanks alot!!

OpenStudy (anonymous):

ur welcome.

OpenStudy (anonymous):

Hmm there is no answer in the choice given.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Sorry, what do u say?

OpenStudy (anonymous):

hmm the answer u gave me is not in the choice given. The choice given : (a) y-3 = 2(x-1) (b) y-2=x-1 (c) y-2=5(x-1) (d) y+4=5(x-2) (e) y-4=5(x-2)

OpenStudy (anonymous):

I don't know what to tell u, my equation satisfies f'(2)=3,f'(1)=5 (ie if you substitute the values in, it equates) Let me try subbing the values in the equations above.

OpenStudy (anonymous):

Hmm okay.:)

OpenStudy (anonymous):

None of the equations given in your choices satisfy both of the conditions. Either the question is wrong or the answers are.

OpenStudy (anonymous):

Yea.thats the problem and there is no suggested answer for it.

OpenStudy (anonymous):

Are u sure both of these are right? f'(2)=3,f'(1)=5

OpenStudy (anonymous):

Note the dash after the f.

OpenStudy (anonymous):

Yes.double checked.

OpenStudy (anonymous):

k, let me check my working again...

OpenStudy (anonymous):

Okay.Thanks again!!

OpenStudy (anonymous):

I stand by my answer if that is the question. What I don't understand is why they have given you the extra pieces of information that u don't need for anything (apparently).

OpenStudy (anonymous):

yes i am kind of condused with the question too.thanks btw:)

OpenStudy (anonymous):

Where is the question from?

OpenStudy (anonymous):

hmm a maths placement test from a university in hong kong.

OpenStudy (anonymous):

I asked on the main board for someone to come check me...

OpenStudy (anonymous):

okay thanks alot!!

OpenStudy (amistre64):

it would appear that you need to find a function that matches the information? f(1)=2 f'(1)=5 f(2)=-2 f'(2)=3 f(5)=1

OpenStudy (anonymous):

Yes, that's the way I read it, can't see why u need the f info though (Mind u quite a few things I don't see these days:-)

OpenStudy (anonymous):

Well, they asked for a tangent line..

OpenStudy (anonymous):

I assumed it didn't matter, they wanted a line so I used the f' info.

OpenStudy (vishweshshrimali5):

k so can I post the link http://openstudy.com/groups/mathematics/updates/4e2ec4e60b8bd8322c635ff5

OpenStudy (amistre64):

f(x) = ax^2 +bx + c ; f'(x) = 2ax + b f'(1) = 2(1)a + b = 5 f'(2) = 2(2)a + b = 3 2a + b = 5 ; *-1 4a + b = 3 -2a -b = -5 4a +b = 3 ----------- 2a = -2 a = -1 ............................ 4(-1) + b = 3 b = 7 ....................... this look right so far?

OpenStudy (anonymous):

That's what I got.

OpenStudy (amistre64):

good; then lets see where this takes me :) f(1) = -(1)^2 +7(1) +c = 2 -1 + 7 + c = 2 c = -4 f(2) = -(2)^2 + 7(2) + c = -2 -4 + 14 + c = -2 c = -12 ... hmm, try that like this then f(1) = (-1)^2 + 7(1) + c = 2 1 + 7 + c = 2 c = -6 f(2) = (-2)^2 + 7(2) + c = -2 4 + 14 + c = -2 c = -20 .... it might not be quadratic ???

OpenStudy (amistre64):

this might help.... given the points (1,2) (2,-2) and (5,1); the slope at (1,2) = 5 the slope at (2,-2) = 3 the equation of the tangent line is: y-Py = m(x-Px) to wit: y-2 = 5(x-1) y+2 = 3(x-2) from your answer choices tho: (a) y-3 = 2(x-1) (b) y-2=x-1 (c) y-2=5(x-1) <-- this matches (d) y+4=5(x-2) (e) y-4=5(x-2)

OpenStudy (anonymous):

Hmm how do you get the slope at (1,2) and (2,-2)?

OpenStudy (anonymous):

How does c match f'(2) = 3?

OpenStudy (amistre64):

the question is simply asking for which answer agrees with any of the information presented. f'(x) IS slope, so we have to use the information given and adapt it to fit the answers

OpenStudy (anonymous):

Or even f'(1) = 5

OpenStudy (anonymous):

f'x is not slope...

OpenStudy (amistre64):

f'(x) defines the slope of the line at a given point ...

OpenStudy (anonymous):

f'x is a line

OpenStudy (amistre64):

f'(1) = 5 means that the slope of the tangent line at f(1) IS 5

OpenStudy (anonymous):

I read it as tangent line with slope as derivative...duh!

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

thanks alot!!

OpenStudy (amistre64):

youre welcome; once I saw the answer choices it made more sense of what we needed to do ...